OpenStudy (anonymous):

A coin was flipped 30 times. Side Frequency Heads 16 Tails 14 Is the experimental or theoretical probability higher for the coin landing heads side up?

OpenStudy (anonymous):

yes because the amount achieved is greater

OpenStudy (anonymous):

? to which was that ment

OpenStudy (anonymous):

heads

OpenStudy (anonymous):

experimental probability = 16/30 theoretical probability=15/30 You tell me which one is greater :)

OpenStudy (anonymous):

its not asking that its asking which theory it is e or t

OpenStudy (anonymous):

HEADS lol

OpenStudy (anonymous):

i think its e

OpenStudy (anonymous):

experimental probability because 16/30 is greater than 15/30 which is the theoritical probability

OpenStudy (anonymous):

ahh i spelled theoretical wrong

OpenStudy (anonymous):

yah it is experamental probabalility